OpenStudy (anonymous):

A ferris wheel (whose seats are 20 m from the axis of rotation) starts from rest and acquires a rotational velocity of 0.5 rad/s completing one revolution. Determine the Ferris wheel's rotational acceleration (assumed to be constant).

OpenStudy (anonymous):

you need centripetal acceleration \[a _{c}=v ^{2}/r\] also you'll probably need to change rads to meters i your velocity 1 rad gives an arc length equal to the radius so 0.5rad/s = 10m/s then \[a _{c}= 10^{2}/20 = 100/20= 5m/s ^{2}\]